OpenStudy (anonymous):

SOmeone help me please 2. The half-life of a radioactive substance is the time it takes for half of the substance to decay. Phosphorus-32 has a half-life of 14.3 days. a. Write an exponential decay function for a 50 mg sample. b. How much phosphorus remains after 84 days?

OpenStudy (anonymous):

half life means you can use \(\frac{1}{2}\) as the base of your exponential function

OpenStudy (anonymous):

starting with 50 grams means the 50 goes out front

OpenStudy (anonymous):

14.3 days means divide \(t\) by \(14.3\)

OpenStudy (anonymous):

set it up as \[\huge f(t)=50\left(\frac{1}{2}\right)^{\frac{t}{14.3}}\]

OpenStudy (anonymous):

f(t)=50

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

i wrote \(f(t)\) above

OpenStudy (anonymous):

i know?

OpenStudy (anonymous):

im lost :/